# references.bib

### v3

@@ 2.1 Baseline methods
If we have a remote sensing data product with a large number $N$ of predictions $\hat{Y}_{1},\hat{Y}_{2},\dots\hat{Y}_{N}$ over the area of interest, we could treat these predictions like ground truth points and use the the formula above with $\hat{Y}$ in place of $Y$. This would produce a much smaller confidence interval, since we would use $N$ instead of $n$ in the denominator of the standard error. However, this map-only estimator approach may result in a biased estimate, and the confidence interval is not guaranteed to have a 95% coverage rate of the true parameter of interest. This is because the map-only estimator does not account for error or uncertainty in the model used to generate the map product predictions.
2.2 Stratified estimator
@@
We first use a map product to partition the area of interest into $K$ strata (sub-areas) such that the target variable is predicted to have low variance within each stratum. Then we sample $n_{k}$ ground truth points uniformly at random from each stratum $1\leq k\leq K$. We can choose $n_{k}$ to be larger for strata predicted to have higher uncertainty to decrease the standard error within the strata and hence make the overall standard error smaller with the same total number of samples.
Suppose we again wish to estimate $\theta^{\star}=E[Y]$ using a stratified random sample. For each stratum $k$, we compute the stratum sample mean $\hat{\theta}_{k}$ and stratum standard error $SE_{k}$ using the same formulae as the classical method. Suppose that in the map product, each stratum $k$ takes up proportion $10<A_{k}<$ 1 of the total area. Then, we compute the stratified mean
@@
$SE_{\text{strat}}=\sqrt{\sum_{k=1}^{K}A_{k}^{2}\cdot SE_{k}^{2}}$
@@
$C_{\text{strat}}=(\hat{\theta}_{\text{strat}}-1.96\cdot SE_{\text{strat}},\ \hat{\theta}_{\text{strat}}+1.96\cdot SE_{\text{strat}}).$
@@
The advantage of stratified estimation with an accurate map and well-chosen sampling strategy is that the variance within each stratum will be small, which results in a smaller overall confidence interval compared to the classical method. If the map predictions are inaccurate, then the ground truth points likely have higher variance within each map stratum, resulting in a larger confidence interval. Thus the uncertainty quantification incorporates the accuracy of the map product predictions.
@@
For example, in our Brazilian Amazon deforestation area estimation use case, we use a remotely sensed stratification map created by [bullock2020satellite] using Landsat surface reflectance imagery at 30 m resolution. The map (Figure 3) partitions the Amazon into seven strata related to forest change from 1995-2017: stable forest, stable non-forest, deforestation, natural disturbance, natural disturbance and deforestation, regrowth, and buffer. The strata are designed so that points within each stratum have low variance in deforestation status (for instance, most points in the “stable forest” and “stable non-forest” strata do not experience deforestation in the time period of interest). As a result, our stratified confidence interval for deforestation area is smaller than the classical confidence interval.
@@
Figure 3: Stratified estimation is a standard approach with remote sensing maps. This example shows a map of the Brazilian Amazon used for stratification. The strata are designed so that points within each stratum have low variance in deforestation status, which reduces the uncertainty when estimating area deforested. Adapted from [bullock2020satellite].
2.3 Prediction-Powered Inference
Prediction-powered inference (PPI) [angelopoulos2023prediction] is an uncertainty quantification method that computes confidence intervals using a small set of ground truth points and a much larger set of machine learning generated predictions. While stratified sampling requires us to commit to a map product and stratification scheme before sampling the ground truth points, the map product used in PPI can be chosen after sampling. The main idea is that PPI assesses the bias in the map product using the ground truth data and then corrects it.
PPI evaluates the accuracy of the map product by comparing the ground truth labels $Y_{1},Y_{2},\dots Y_{n}$ to the map product predictions $\hat{Y}_{1},\hat{Y}_{2},\dots\hat{Y}_{n}$ at the corresponding ground truth point locations. For a concrete example, suppose we wish to estimate $\theta^{\star}=E[Y]$ as before. The PPI point estimate is
$\hat{\theta}_{\text{PPI}}=\underbrace{\frac{1}{N}\sum_{i=1}^{N}\hat{Y}_{i}^{\prime}}_{\text{map-only estimate}}-\underbrace{\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}-Y_{i})}_{\text{bias correction}}.$
The PPI standard error is
$SE_{\text{PPI}}$ $=\sqrt{\frac{1}{N}\mathrm{Var}(\hat{Y}_{i}^{\prime})+\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})}$
$\approx\sqrt{\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})},$
where the approximation holds when $N$ is much larger than $n$. This yields the PPI 95% confidence interval,
$C_{\text{PPI}}=(\hat{\theta}_{\text{PPI}}-1.96\cdot SE_{\text{PPI}},\ \hat{\theta}_{\text{PPI}}+1.96\cdot SE_{\text{PPI}}).$
The first term in $SE_{\text{PPI}}$ is the variance corresponding to using only map predictions to estimate $\theta$, i.e. $\frac{1}{N}\mathrm{Var}(\hat{Y}_{i}^{\prime})$. The second term is the variance from the bias correction, $\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})$. As the number of map product points $N$ grows large, $SE_{\text{PPI}}$ approaches $\sqrt{\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})}$. This will be a realistic approximation in many cases, since remote sensing maps at a global, national, or province level can contain millions or even billions of pixels, whereas the number of ground truth points $n$ is commonly in the hundreds or thousands. The greater the accuracy of the map product, the smaller $\mathrm{Var}(\hat{Y}_{i}-Y_{i})$ will be and the smaller the PPI confidence interval will be. Note that the PPI confidence interval is smaller than the classical confidence interval if $\mathrm{Var}(\hat{Y}_{i}-Y_{i})<\mathrm{Var}(Y_{i})$, which will be true for any map whose predictions correlate positively with the ground truth.
PPI can also give confidence intervals for linear and logistic regression coefficients, but the formulae are more complex. We sketch the main idea here and refer the reader to [angelopoulos2023ppi++, angelopoulos2023prediction] for a full exposition. In brief, PPI begins like a map-only estimator by using a large set of map predictions $\hat{Y}^{\prime}_{1},\hat{Y}^{\prime}_{2},\dots\hat{Y}^{\prime}_{N}$ to compute a measure of fit $m_{\theta}$. Then, PPI computes a bias correction $\Delta_{\theta}$ by comparing the ground truth data $Y_{i}$ to the predictions $\hat{Y}_{i}$. The relevant formula for the measure of fit and bias correction depends on what we seek to estimate — areas, binary regression coefficients, linear regression coefficients, etc. — see Table 3. Lastly, the map-only $m_{\theta}$ and the bias correction $\Delta_{\theta}$ are combined to construct the PPI 95% confidence interval
$C_{\text{PPI}}=\{\theta:|m_{\theta}+\Delta_{\theta}|\leq w_{\theta}\}$
where $w_{\theta}$ is a constant that depends on the confidence level (95% in this case).
@@ 2.3 Prediction-Powered Inference
PPI can be used with either uniform random sampling or stratified random sampling of the ground truth points [fisch2024stratified]. Stratified (weighted) PPI can be applied as follows for mean estimation. Each ground truth value $Y_{i}$ and labeled prediction $\hat{Y}_{i}$ is multiplied by weight $w_{i}=A_{k}\frac{n}{n_{k}}$ where $k$ is the stratum where point $i$ is located, $A_{k}$ is the area proportion of stratum $k$ and $n_{k}$ is the number of ground truth samples in stratum $k$. Similarly, each unlabeled prediction $\hat{Y}^{\prime}_{i}$ in stratum $k$ is multiplied by weight $w_{i}=A_{k}\frac{N}{N_{k}}$ where $N_{k}$ is the number of map product points in stratum $k$. Then the same PPI formulae as above can be used.
Estimand Measure of fit $m_{\theta}$ Bias correction $\Delta_{\theta}$
Mean $\theta-\frac{1}{N}\sum_{i=1}^{N}\hat{Y}_{i}^{\prime}$ $\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}-Y_{i})$
Logistic regression $\frac{1}{N}\sum_{i=1}^{N}\left(\frac{X_{i}^{\prime}}{1+e^{-\theta^{T}X_{i}^{\prime}}}-\hat{Y}_{i}^{\prime}\right)$ $\frac{1}{n}\sum_{i=1}^{n}X_{i}\left(\hat{Y}_{i}-Y_{i}\right)$
Linear regression $\theta-(X^{\prime})^{\dagger}\hat{Y}^{\prime}$ $X^{\dagger}\left(\hat{Y}-Y\right)$
Table 3: Prediction-powered inference (PPI) formulae for three common estimation tasks found in remote sensing: mean estimation, logistic regression coefficients, and linear regression coefficients.
We use the ppi-python software package to implement the unweighted and weighted versions of PPI for our use cases. The software includes a tuning parameter $\lambda$ that we set using the map product points and ground-truth points. The confidence intervals are unbiased even with this data-driven tuning [angelopoulos2023ppi++].
2.4 Post-stratified area estimator
@@
The post-stratified area estimator, commonly used in the remote sensing community, utilizes the confusion matrix between ground truth and map product land cover labels to estimate the area of a land cover class [card1982using, stehman2013estimating, olofsson2013making]. Like PPI, the post-stratified area estimator computes confidence intervals by using ground truth points to correct for bias in the map product estimates. This can be viewed as a correction that uses the map errors relative to the ground truth points to modify the map-only area estimates. We show mathematically in Appendix A that PPI and the post-stratified area estimator result in similar confidence intervals as the number of map product points $N$ grows large. However, unlike PPI, the post-stratified area estimator only applies to area estimation but not regression coefficient estimation.
@@
Turning to the method, suppose we have a map product with $K$ land cover classes, as well as $n$ randomly sampled ground truth points. The post-stratified area estimator constructs the $K\times K$ misclassification confusion matrix between ground truth points and map product predictions at the ground truth point locations. For each $1\leq i\leq K$ and $1\leq j\leq K$, let $n_{ij}$ denote the number of points that are in map class $i$ and ground truth class $j$. For each class $1\leq i\leq K$, let $10<A_{i}<$ 1 denote the proportion of the total map area taken up by the map class.
@@
Then, the post-stratified area estimate for the area of class $j$ is
@@
$SE_{j,\text{post}}=\sqrt{\sum_{i=1}^{K}A_{i}^{2}\frac{\frac{n_{ij}}{n_{i\cdot}}(1-\frac{n_{ij}}{n_{i\cdot}})}{n_{i\cdot}}}.$
The 95% confidence interval for the area of class $j$ is
$C_{j,\text{post}}=(\hat{\theta}_{j,\text{post}}-1.96\cdot SE_{j,\text{post}},\ \hat{\theta}_{j,\text{post}}+1.96\cdot SE_{j,\text{post}}).$
Area estimation is a special case of mean estimation in which we estimate the mean of a binary variable that is $1$ for all points in class $j$ and 00 for all points outside of class $j$. Thus, we can also perform area estimation using methods such as PPI and stratified estimation. If the $n$ ground truth points are obtained by stratified random sampling using the $K$ map land cover classes as the strata, the post-stratified confidence interval is identical to the stratified confidence interval.
2.5 Evaluating performance using effective sample size
We use effective sample size to compare the effectiveness of uncertainty quantification methods. Suppose a classical confidence interval of width $w$ is generated using $n$ ground truth points. If another uncertainty quantification method uses $n$ ground truth points (along with a map product) to generate a confidence interval of width $w^{\prime}$, then the effective sample size of the new method is
$n_{\text{effective}}=n\cdot\left(\frac{w}{w^{\prime}}\right)^{2}.$
This is the number of ground truth points that would be required to generate a classical confidence interval of width $w^{\prime}$.
3 Area estimation
In this section, we demonstrate uncertainty quantification for two area estimation use cases. First, we simulate estimating maize area in Iowa in 2022, artificially degrading the map product quality relative to simulated “ground truth” points, in order to observe the effect of map product noise and bias on the area estimate under different methods. We will see through this simulation the importance of bias correction before using maps to estimate total area. We then use real ground truth data and real map products developed by researchers to estimate change in tree cover in the Brazilian Amazon in 2000–2015. We compare the classical, map-only (pixel-counting), stratified, post-stratified, PPI, and stratified PPI estimators.

### v4

@@ Appendix D PPI for mean estimation
If we have a remote sensing data product with a large number $N$ of predictions $\hat{Y}_{1},\hat{Y}_{2},\dots\hat{Y}_{N}$ over the area of interest, we could treat these predictions like ground truth points and use the the formula above with $\hat{Y}$ in place of $Y$. This would produce a much smaller confidence interval, since we would use $N\gg n$ in the denominator of the standard error. However, this map-only estimator approach may result in a biased estimate, and the confidence interval is not guaranteed to have a 95% coverage rate of the true parameter of interest. This is because the map-only estimator does not account for error or uncertainty in the model used to generate the map product predictions.
Prediction-Powered Inference: Map + ground truth.
PPI [angelopoulos2023prediction] evaluates the accuracy of the map product by comparing the ground truth labels $Y_{1},Y_{2},\dots,Y_{n}$ to the map product predictions $\hat{Y}_{1},\hat{Y}_{2},\dots,\hat{Y}_{n}$ at the corresponding ground truth point locations. This calibration set of $n$ points is a simple or stratified random sample over the region of interest, and it must be separate from the training dataset used to train the machine learning model. PPI uses the calibration set to correct for bias in the map-only estimate computed from the $N$ unlabeled map points $\hat{Y}^{\prime}_{1},\hat{Y}^{\prime}_{2},\dots\hat{Y}^{\prime}_{N}$.
For a concrete example, suppose we wish to estimate $\theta^{\star}=E[Y]$ as before. In the special case of estimating $\theta^{\star}=E[Y]$, the PPI point estimate is
$\hat{\theta}_{\text{PPI}}=\underbrace{\frac{1}{N}\sum_{i=1}^{N}\hat{Y}_{i}^{\prime}}_{\text{map-only estimate}}-\underbrace{\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}-Y_{i})}_{\text{bias correction}},$
which coincides precisely with an estimator previously studied in the remote sensing literature (see Equation 4c of [MCROBERTS2022113168]). Assuming the map sample and calibration sample are independent, the standard deviation of this estimator is
$\sigma_{\text{PPI}}$ $=\sqrt{\frac{1}{N}\mathrm{Var}(\hat{Y}_{i}^{\prime})+\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})}$
$\approx\sqrt{\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})},$
where the approximation holds when $N$ is much larger than $n$. Letting $\hat{\sigma}_{\text{PPI}}$ be an estimate of $\sigma_{\text{PPI}}$ based on plugging sample variance estimates into the formula for $\sigma_{\text{PPI}}$, the PPI 95% confidence interval is given by
$C_{\text{PPI}}=(\hat{\theta}_{\text{PPI}}-1.96\cdot\hat{\sigma}_{\text{PPI}},\ \hat{\theta}_{\text{PPI}}+1.96\cdot\hat{\sigma}_{\text{PPI}}).$
The first term in $\sigma_{\text{PPI}}$ is the variance corresponding to using only map predictions to estimate $\theta$, i.e. $\frac{1}{N}\mathrm{Var}(\hat{Y}_{i}^{\prime})$. The second term is the variance from the bias correction, $\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})$. As the number of map product points $N$ grows large, $\sigma_{\text{PPI}}$ approaches $\sqrt{\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})}$. This will be a realistic approximation in many cases, since remote sensing maps at a global, national, or province level can contain millions or even billions of pixels, whereas the number of ground truth points $n$ is commonly in the hundreds or thousands. The greater the accuracy of the map product, the smaller $\mathrm{Var}(\hat{Y}_{i}-Y_{i})$ will be and the smaller the PPI confidence interval will be. Note that the PPI confidence interval is smaller than the GT-only confidence interval if $\mathrm{Var}(\hat{Y}_{i}-Y_{i})<\mathrm{Var}(Y_{i})$.
Like PPI, the post-stratified area estimator combines a map product with a small amount of ground truth data to make a precise and unbiased estimate. The post-stratified estimator and PPI produce similar outputs for area estimation; for more detailed comparisons of the two methods, see Appendices E and F. However, the post-stratified estimator only applies to area estimation, not regression coefficient estimation, while PPI can be applied to both (Figure 1 in the main text). Similarly, an estimator from [MCROBERTS2022113168] coincides precisely with the PPI estimator in the context of mean estimation, but does not apply to the setting of regression coefficient estimation.
Appendix E Area estimation experiments
In this appendix, we apply PPI and other methods to area estimation tasks with both ground truth and remotely sensed data.
We first describe two well-known methods for area estimation (stratified estimator and post-stratified estimator). These methods were introduced in [cochran1977sampling] and are commonly used in the remote sensing community [olofsson2013making, olofsson2014good]. PPI for area estimation (described in Appendix D) produces similar confidence intervals as the post-stratified estimator; for a mathematical comparison, see Appendix F.
Then, we apply PPI, the stratified estimator, and the post-stratified estimator to two area estimation use cases (Iowa maize crop area and Brazilian Amazon deforestation area). We also compare results with the GT-only and map-only mean estimation methods described in Appendix D.
E.1 Stratified estimator
@@
We first use a map product to partition the area of interest into $K$ strata (sub-areas) such that the response variable is predicted to have low variance within each stratum. Then we sample $n_{k}$ ground truth points uniformly at random from each stratum $1\leq k\leq K$. We can choose $n_{k}$ to be larger for strata predicted to have higher uncertainty to decrease the standard error within the strata and hence make the overall standard error smaller with the same total number of sample units.
Suppose we observe the response variable $Y$ over a given area, and we wish to estimate $\theta^{\star}=E[Y]$ using a stratified random sample. For each stratum $k$, we compute the stratum sample mean $\hat{\theta}_{k}$ and stratum standard error $\hat{\sigma}_{k}$ using the same formulae as the GT-only method from Appendix D. Suppose that in the map product, each stratum $k$ takes up proportion $10<A_{k}<$ 1 of the total area (with $\sum_{k=1}^{K}A_{k}=1$). Then, we compute the stratified mean
@@
$\hat{\sigma}_{\text{strat}}=\sqrt{\sum_{k=1}^{K}A_{k}^{2}\cdot\hat{\sigma}_{k}^{2}}$
@@
$C_{\text{strat}}=(\hat{\theta}_{\text{strat}}-1.96\cdot\hat{\sigma}_{\text{strat}},\ \hat{\theta}_{\text{strat}}+1.96\cdot\hat{\sigma}_{\text{strat}}).$
@@
The advantage of stratified estimation with an accurate map and well-chosen sampling strategy is that the variance within each stratum will be small, which results in a smaller overall confidence interval compared to the GT-only method. If the map predictions are inaccurate, then the ground truth points likely have higher variance within each map stratum, resulting in a larger confidence interval. Thus the uncertainty quantification incorporates the accuracy of the map product predictions.
@@
For example, in our Brazilian Amazon deforestation area estimation use case, we use a remotely sensed stratification map created by [bullock2020satellite] using Landsat surface reflectance imagery at 30 m resolution. The map (Figure 7) partitions the Amazon into seven strata related to forest change from 1995-2017: stable forest, stable non-forest, deforestation, natural disturbance, natural disturbance and deforestation, regrowth, and buffer. The strata are designed so that points within each stratum have low variance in deforestation status (for instance, most points in the “stable forest” and “stable non-forest” strata do not experience deforestation in the time period of interest). As a result, our stratified confidence interval for deforestation area is smaller than the GT-only confidence interval.
@@
Figure 7: Stratified estimation is a commonly used approach with remote sensing maps. This example shows a map of the Brazilian Amazon used for stratification. The strata are designed so that points within each stratum have low variance in deforestation status, which reduces the uncertainty when estimating area deforested. Adapted from [bullock2020satellite].
@@
PPI can be used with either uniform random sampling or stratified random sampling of the ground truth points [fisch2024stratified]. Stratified (weighted) PPI can be applied as follows for mean estimation. Each ground truth value $Y_{i}$ and labeled prediction $\hat{Y}_{i}$ is multiplied by weight $w_{i}=A_{k}\frac{n}{n_{k}}$ where $k$ is the stratum where point $i$ is located, $A_{k}$ is the area proportion of stratum $k$, $n_{k}$ is the number of ground truth sample units in stratum $k$, and $n=\sum_{i=1}^{K}n_{k}$. (Note that the weights are designed such that the sum of weights of all $n_{k}$ ground truth points in stratum $k$ is $A_{k}n$, and the sum of weights of all $n$ ground truth points is $\sum_{i=1}^{K}A_{k}n=n$. If the points were uniformly randomly sampled, this would be equivalent to only having a single “stratum,” and each point would have weight 1.) Similarly, each unlabeled prediction $\hat{Y}^{\prime}_{i}$ in stratum $k$ is multiplied by weight $w_{i}=A_{k}\frac{N}{N_{k}}$ where $N_{k}$ is the number of map product points in stratum $k$ and $N=\sum_{i=1}^{K}N_{k}$. Then the same PPI formulae as in Appendix D can be used.
E.2 Post-stratified area estimator
@@
The post-stratified area estimator, commonly used in the remote sensing community, utilizes the confusion matrix between ground truth and map product land cover labels to estimate the area of a land cover class [card1982using, stehman2013estimating, olofsson2013making]. Like PPI, the post-stratified area estimator computes confidence intervals by using ground truth points to correct for bias in the map product estimates. This can be viewed as a correction that uses the map errors relative to the ground truth points to modify the map-only area estimates. We show mathematically in Appendix F that PPI and the post-stratified area estimator result in similar confidence intervals as the number of map product points $N$ grows large. However, unlike PPI, the post-stratified area estimator only applies to area estimation but not regression coefficient estimation.
@@
Turning to the method, suppose we have a map product with $K$ land cover classes, as well as $n$ randomly sampled ground truth points. The post-stratified area estimator constructs the $K\times K$ misclassification confusion matrix between ground truth points and map product predictions at the ground truth point locations. For each $1\leq i\leq K$ and $1\leq j\leq K$, let $n_{ij}$ denote the number of points that are in map class $i$ and ground truth class $j$. Also let $n_{i\cdot}=\sum_{j=1}^{K}n_{ij}$ denote the total number of ground truth samples that have map class $i$. For each class $1\leq i\leq K$, let $10<A_{i}<$ 1 denote the proportion of the total map area taken up by the map class.
@@
Then, motivated by the formula $\mathbb{P}(Y=j)=\sum_{i=1}^{K}\mathbb{P}(\hat{Y}=i)\cdot\mathbb{P}(Y=j\ |\ \hat{Y}=i)$, the post-stratified area estimate for the proportion of area of class $j$ is
@@
$\hat{\sigma}_{j,\text{post}}=\sqrt{\sum_{i=1}^{K}A_{i}^{2}\frac{\frac{n_{ij}}{n_{i\cdot}}(1-\frac{n_{ij}}{n_{i\cdot}})}{n_{i\cdot}}}.$
The 95% confidence interval for the proportion of area of class $j$ is
$C_{j,\text{post}}=(\hat{\theta}_{j,\text{post}}-1.96\cdot\hat{\sigma}_{j,\text{post}},\ \hat{\theta}_{j,\text{post}}+1.96\cdot\hat{\sigma}_{j,\text{post}}).$
Area estimation is a special case of mean estimation in which we estimate the mean of a binary variable that is $1$ for all points in class $j$ and 00 for all points outside of class $j$. Thus, we can also perform area estimation using methods such as PPI and stratified estimation.
If the $n$ ground truth points are obtained by stratified random sampling using the $K$ map land cover classes as the strata, the post-stratified confidence interval computed using this data is identical to the stratified confidence interval [olofsson2013making]. However, unlike the stratified estimator, the post-stratified estimator can be applied even if the ground truth points are obtained using uniform random sampling.
E.3 Area estimation examples
In this section, we demonstrate uncertainty quantification for two area estimation use cases. First, we simulate estimating maize area in Iowa in 2022, artificially degrading the map product quality relative to simulated “ground truth” points, in order to observe the effect of map product noise and bias on the area estimate under different methods. We will see through this simulation the importance of bias correction before using maps to estimate total area. We then use real ground truth data and real map products developed by researchers to estimate change in tree cover in the Brazilian Amazon in 2000–2015. We compare the GT-only, map-only (pixel-counting), stratified, post-stratified, PPI, and stratified PPI estimators.